OpenStudy (anonymous):

Given vectors u= 4i+9j and v= -3i+ 7j, find 7u-7v in terms of i and j

OpenStudy (zzr0ck3r):

7u = 7*4i+9*7j -7v = (-7)(-3)i+(-7)(7)j 7u-7v=7u+(-7v) = 7*4i+9*7j+(-7)(-3)i+(-7)(7)j=7*4i+(-7)(-3)i+9*7j+(-7)(7)j = (7*4+(-7)(-3))i+(9*7+(-7)(7))j

OpenStudy (anonymous):

Thank you so much!

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

I prob coult have done the easy arithmetic and made it look better 7u=21i-49j ....

OpenStudy (anonymous):

The answer options are A: 9i+48j B: 52i+48j C: 15i+48j D: 49i+14j